OpenStudy (anonymous):

log(b)√3x=_____ write as a sum of multiples of logarithms.

OpenStudy (anonymous):

now go the other way \[\log(\sqrt{3x})=\frac{1}{2}\log(3x)=\frac{1}{2}(\log(3)+\log(x))\]

OpenStudy (anonymous):

\[log_b\sqrt{3x}=\frac{1}{2}log_b(3x)\] \[\frac{1}{2}log_b(3x)=\frac{1}{2}(log_b3+log_bx)\]